### v3

# Beyond KAN: Introducing KarSein for Adaptive High-Order Feature Interaction Modeling in CTR Prediction

Abstract
Modeling feature interactions is crucial for click-through rate (CTR) prediction, particularly when it comes to high-order explicit interactions. Traditional methods struggle with this task because they often pre-define a maximum interaction order, which relies heavily on prior knowledge and can limit the model’s effectiveness. Additionally, modeling high-order interactions typically leads to increased computational costs. Therefore, the challenge lies in adaptively modeling high-order feature interactions while maintaining efficiency. To address this issue, we introduce Kolmogorov-Arnold Represented Sparse Efficient Interaction Network (KarSein), designed to optimize both predictive accuracy and computational efficiency. We firstly identify limitations of directly applying Kolmogorov-Arnold Networks (KAN) to CTR, and then introduce KarSein to overcome these issues. It features a novel architecture that reduces the computational costs of KAN and supports embedding vectors as feature inputs. Additionally, KarSein employs guided symbolic regression to address the challenge of KAN in spontaneously learning multiplicative relationships. Extensive experiments demonstrate KarSein’s superior performance, achieving significant predictive accuracy with minimal computational overhead. Furthermore, KarSein maintains strong global explainality while enabling the removal of redundant features, resulting in a sparse network structure. These advantages also position KarSein as a promising method for efficient inference  [^1] https://github.com/Ancientshi/KarSein.
Introduction
In the ever-evolving realm of digital advertising and recommendation systems, Click-Through Rate (CTR) prediction has emerged as a pivotal element for optimizing user engagement and enhancing revenue streams. Despite its seemingly straightforward goal, CTR prediction methods aim at capturing underlying feature relationships among the complicated context data.
Figure 1: Comparison between KarSein and pioneer methods. KarSein merits in building interactions without space transformation, capturing high order features simply via activation, it provides global explanations so can remove redundant features.
Typically, these methods either capture implicit feature interaction by directly modeling the fusion of all features using deep neural networks (Covington, Adams, and Sargin 2016; Cheng et al. 2016), or learn explicit feature interaction by manually defining the interaction form or order through factorization-based models (Covington, Adams, and Sargin 2016; Cheng et al. 2016). Due to the exponential growth of combinational complexity, explicit learning methods typically limit themselves to modeling low-order interactions and only integer-order interactions, which hinders their ability to accurately represent scenarios requiring high-order and nuanced feature interactions. Most recently, AFN (Cheng, Shen, and Huang 2020) and EulerNet (Tian et al. 2023) design novel feature transformations to adaptively learn high-order interactions. However, they require a preceding embedding space and face challenges such as information loss, numerical stability issues, and computational overhead during transformations. Therefore, here is the first challenge: how to develop a method that can effectively model high-order feature interactions in different real-world scenarios? Moreover, existing CTR researches (Xiao et al. 2017; Weiping et al. 2018; Li et al. 2019; Huang, Zhang, and Zhang 2019) combine the high-order features relying on the specific context, leading to unsufficient and localized explainability. The absence of convincing rationales behind model predictions calls their reliability and security into question. In various applications, such as medication recommendation and financial services, untrustworthy and unreliable advertisements can lead to severe consequences, including economic loss or health issues. So the second challenge is: how to construct a model with global explanations to improve the trustworthy and decrease the feature redundancy for CTR prediction?
Recently, a novel neural network architecture known as Kolmogorov-Arnold Networks (KAN) (Liu et al. 2024) has been proposed, garnering significant attention within the deep learning community for its exceptional data fitting capabilities, intuitive interpretability, and high structural sparsity. It leads us to contemplate the significant potential of applying KAN in feature interaction modeling for CTR prediction. However, applying vanilla KAN for CTR prediction will result in inferior performance suffering from these problems. (1) The sensitivity to regularization settings and network structure initialization makes it challenging to generalize in CTR prediction, preventing even simple multiplicative relationships between basic features. (2) KAN not supports 2D embedding vectors as input features, limits it’s modeling feature interactions at vector-wise manner. (3) The high computational complexity of KAN is impractical for CTR.
To this end, we propose Kolmogorov–Arnold Represented Sparse Efficient Interaction Network, named as KarSein, for CTR prediction. Unlike KAN, which employs multiple activation functions per input feature, KarSein is designed to allocate just one activation function per feature, thereby significantly streamlining computations. Additionally, we extend KAN to support two-dimensional input features, allowing embedding vectors to serve as input features for modeling feature interactions at vector-wise level. Furthermore, in each KarSein layer, we incorporate an optional step of pairwise multiplication between layer’s input features and basic features for enforcing the network to learn multiplicative relationships, this will further facilitating latter layers capturing more multiplicative interactions and contribute to prediction greatly. KarSein also retains the simplification technology of KAN, the strong global explainability can support feature “de-redundancy” for all instances.
In summary, our paper offers the following contributions:
  • We are the first to explore the application of KAN in CTR prediction, identifying critical limitations and providing insightful findings.
  • We introduce the Kolmogorov–Arnold Represented Sparse Efficient Interaction Network (KarSein), an innovative CTR prediction model that represents a novel approach in the field.
  • We demonstrate our proposed KarSein achieves state-of-the-art performance and lowest computational cost across three datasets.
  • We point out KarSein’s potential for structural sparsity learning, which could further significantly accelerate model inference efficiency.
Preliminaries
Problem Formulation for CTR
Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. For a user-item pair $(u,i)\in\mathcal{U}\times\mathcal{I}$, we define $\mathbf{x}_{u,i}=[x_{1},\ldots,x_{m}]$ as the feature vector capturing relevant attributes, including categorical (e.g., user and item IDs) and numerical (e.g., age) features. The Click-Through Rate (CTR) prediction task aims to estimate $P(y=1\mid\mathbf{x}_{u,i})$, where $y\in\{0,1\}$ indicates whether the user clicked on the item. Formally, we define $\hat{y}=f(\mathbf{x}_{u,i};\Theta)$ where $\hat{y}$ is the predicted click probability, $f$ is the predictive model, and $\Theta$ represents the model parameters. Given a training dataset $\mathcal{D}=\{(u_{j},i_{j},\mathbf{x}_{(u_{j},i_{j})},y_{j})\}_{j=1}^{N}$ of $N$ instances, we optimize $\Theta$ to minimize the prediction error. A common approach is to minimize the log loss: $\mathcal{L}(\Theta)=-\frac{1}{N}\sum_{j=1}^{N}\left[y_{j}\log\hat{y}_{j}+(1-y_{j})\log(1-\hat{y}_{j})\right]$
Feature Interactions Modeling
In recommendation systems, the model typically includes a trainable parameterized embedding layer, denoted as $E(\cdot)$. Given a categorical ID feature $x_{j}$, the model first maps $x_{j}$ into dense vectors in a low-dimensional latent space: $e_{j}=E(x_{j})\in\mathbb{R}^{D}$, where $D$ is the dimension of the embedding. Feature interaction modeling is then conducted on these vectorized embeddings. Feature interaction types can be broadly categorized into two primary paradigms: implicit and explicit. They often complement with each other, and more explanatory can be seen in Appendix D.
Explicit Feature Interactions
These interactions typically occur via multiplication at the vector level among representing vectors of basic field features. For a set of features $\{x_{1},x_{2},\ldots,x_{m}\}$ with corresponding embeddings $\{e_{1},e_{2},\ldots,e_{m}\}$, and consider hadamard product (element-wise multiplication) as the multiplication method, then for the $k$-th order feature interactions, we can enumerate all combinations as $\{e_{n_{1}}\odot e_{n_{2}}\odot\cdots\odot e_{n_{k}}\mid n_{1},n_{2},\ldots,n_{k}\in\{1,2,\ldots,m\}\}$. By stacking all the elements, we form a matrix $\mathbf{X}_{k}\in\mathbb{R}^{m^{k}\times D}$. Concatenating all interactions from the first-order to the $k$-th order results in a matrix $\mathbf{X}_{1\sim k}\in\mathbb{R}^{\ \sum_{i=1}^{k}m^{i}\;\times\;D}$, where each row represents a high order explicit feature interaction.
When constructing high-order features with large $k$, the number of rows in $\mathbf{X}_{1\sim k}$ is given by $s=\sum_{i=1}^{k}m^{i}$, which increases exponentially with $k$, specifically, $s\sim O(m^{k})$. Traditional methods typically construct $\mathbf{X}_{1\sim k}$ with pre-defined $k$ (e.g., $k=2$). Advance methods focus on adaptively learning arbitrary high-order feature interactions without pre-defining $k$ or exhaustively enumerating all features.
Implicit Feature Interactions
Implicit feature interactions refer to the interactions between features that are not explicitly predefined. Instead, these interactions are captured automatically by models. Let $\mathbf{e}=e_{1}||e_{2}||\ldots||e_{m}$ denote the result of wide concatenated embedding vectors. Then let $\mathbf{W}$ be the linear transformation matrix applied to $\mathbf{e}$, and $\sigma$ be the activation function for non-linearities. For an $L$-layer DNN, this implicit feature interaction modeling can be expressed as:
$\text{MLP}(\mathbf{e})=(\mathbf{W}_{L-1}\circ\sigma\circ\mathbf{W}_{L-2}\circ\sigma\circ\cdots\circ\mathbf{W}_{1}\circ\sigma\circ\mathbf{W}_{0})\mathbf{e}$
Figure 2: Comparison of KarSein and KAN Models for CTR Prediction: Network Architecture and Feature Interaction Modeling Process. Both KarSein-implicit (1-right) and KAN (2) models are configured with a 3-3-1 network architecture, while the KarSein-explicit (1-left) model is set with a 3-2-1 architecture. The output of each layer is depicted using a pink-to-orange gradient, indicating the increasing order of feature interactions from low to high.
Implicit feature interactions often involve learning the interaction patterns at the bit level. This process is supported by the Universal Approximation Theorem, which states that a neural network with at least one hidden layer, containing a finite number of neurons, and using an appropriate activation function (such as sigmoid or ReLU), can approximate any complex continuous function.
Kolmogorov-Arnold Network
For a smooth function $f:[0,1]^{m}\rightarrow\mathbb{R}$, the Kolmogorov-Arnold Theorem states:
$f(x)=f(x_{1},\cdots,x_{m})=\sum_{q=1}^{2m+1}\Phi_{q}\left(\sum_{p=1}^{m}\phi_{q,p}(x_{p})\right)$
where $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. This theorem implies that a multivariate function $f(x)$ can be expressed using univariate functions and summation option. This theorem simplifies the task of learning high-dimensional functions by reducing it to learning a polynomial number of one-dimensional functions. However, these one-dimensional functions can be non-smooth or even fractal, making them impractical for direct application in machine learning models. However, Kolmogorov–Arnold Networks (KAN) (Liu et al. 2024) presents an optimistic perspective, it employs highly flexible, learnable B-Spline curves as $\phi(\cdot)$ for activation, and extends the theorem to neural networks with arbitrary widths and depths. Formally, for an activation function $\phi(\cdot)$, let $\text{Silu}(\cdot)$ denote the Silu activation, and $B(\cdot)$ represent the B-Spline curves activation withe grid size $g$ and order $\kappa$. The activation function $\phi(x)$ is defined as $\phi(x)=w_{\phi}(\operatorname{B}(x)+\text{Silu}(x))$, where $\operatorname{B}(x)=\sum_{i=0}^{g+\kappa}c_{i}N_{i,\kappa}(x)$. Here, both $w_{\phi}$ and $c_{i}$ are learnable parameters. Let $\Phi_{L}$ is the activation function matrix corresponding to the L-th KAN layer. A general KAN network is a composition of L layers, given an wide concatenated embedding vector $e$ as input, the output of KAN is:
$\text{KAN}(e)=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{1}\circ\Phi_{0})e.$
KAN for CTR Prediction
Multiplicative Relationship Learning
We find that the Kolmogorov–Arnold Representation Theorem is particularly well-suited for representing feature interactions combined with multiplicative relationship. Given two basic field feature $x_{1}$ and $x_{2}$, the derived second-order feature interaction $x_{1}x_{2}$ can be expressed as $\frac{1}{2}(\Phi_{1}(\phi_{1,1}(x_{1})+\phi_{1,2}(x_{2}))+\Phi_{2}(\phi_{2,1}(x_{1})+\phi_{2,2}(x_{2})))$, where $\Phi_{1}(x)=x^{2}$, $\Phi_{2}(x)=-x$, $\phi_{1,1}(x)=x$, $\phi_{1,2}(x)=x$, $\phi_{2,1}(x)=x^{2}$, $\phi_{2,2}(x)=x^{2}$. Thus, KAN can model feature interactions through symbolic regression, offering greater explainability and efficiency compared to DNN.
We also find that KAN’s symbolic regression ability is sensitive to the structure initialization and the regularization setting, which can see details in Appendix B.1. In more general scenarios, KAN is hard to spontaneously perform right symbolic regression for learning higher-order multiplicative features interactions. This observation underscores a regrettable reality: KAN, akin to DNN, remains intrinsically limited in its ability to autonomously learn multiplicative features interactions. Consequently, relying solely on KAN’s spontaneously learning may not yield optimal CTR prediction results.
Vanilla Application in CTR
We designed a KAN network with layers of width $mD-64-64-1$ trained on the MovieLens-1M dataset, where wide concatenated embedding vectors $e$ serve as the network input. This network outputs a numeric value, which is then activated by sigmoid function for CTR prediction. When model is convergent, achieving an AUC score of 0.8273, whereas a DNN model with the same layer and neuron configuration achieves an AUC of 0.8403. This shows that he vanilla application of KAN for CTR prediction yields suboptimal results. But we find that, in the optimized KAN, many network connections do little contribution, using the network Simplification techniques, the KAN network is pruned from the initial $mD-64-64-1$ neurons down to $mD-1-1-1$ neurons, indicating that a single activation function per feature may be sufficient, eliminating the cost for multiple activation functions on the same feature, which is the parameter-redundancy in KAN. The detailed exploration settings are in Appendix B.2.
KARSEIN
The preliminary exploration of KAN for CTR prediction highlights the limitations of this vanilla methods, but the findings also motivate us to make modifications to adapt KAN for the CTR task. In this section, we introduce KarSein. A schematic of its architecture is shown in Figure 2.
KarSein Interaction Layer
The KarSein model is constituted by multiple stacked KarSein interaction layers. The layer takes a set of embedding vectors as input and generates higher-order feature interactions. The feature interaction consists of three steps on embedding vectors: optional pairwise multiplication, learnable activation transformation, and linear combination.
We begin by describing the KarSein interaction layer at the $L$-th level, where the input dimension is $H_{L-1}$ and the output dimension is $H_{L}$. Here, $L\in[1,\mathcal{T}]$, and $\mathcal{T}$ denotes the depth of the stacked KarSein interaction layers. For the $L$-th layer, let $\mathbf{X}^{L-1}\in\mathbb{R}^{H_{L-1}\times D}$ represent the input matrix, and $\mathbf{X}^{L}\in\mathbb{R}^{H_{L}\times D}$ represent the output matrix. Specifically, we define $H_{0}=m$, and $\mathbf{X}^{0}\in\mathbb{R}^{m\times D}$ as the matrix formed by stacking $e_{1},e_{2},\ldots,e_{m}$. Additionally, we define $H_{\mathcal{T}}=1$ because the final layer is designed to have a single output neuron, ultimately modeling a highly intricate high-order feature interaction, denoted as $\mathbf{X}^{\mathcal{T}}\in\mathbb{R}^{1\times D}$.
Pairwise Multiplication
In our methodology, we conduct feature interactions between $\mathbf{X}^{L-1}$ and $\mathbf{X}^{0}$. This process involves computing pairwise hadamard products of features. We then concatenate the results with $\mathbf{X}^{L-1}$ yielding new matrix for substituting $\mathbf{X}^{L-1}$, which serves as the input for subsequent steps. Generally, incorporating this optional Pairwise Multiplication step only within the first two KarSein interaction layer is adequate to guide the model’s learning can incorporate multiplicative relationships. As shown in Figure 2, the Pairwise Multiplication step generates second-order features $f^{2}(e_{1},e_{2})$. Through high-order activations and linear transformations, these features evolve into more complex multivariate high-order interactions, such as $f^{6}(e_{1},e_{2},e_{3})$, a sixth-degree polynomial feature involving three variables. While $f^{6}(e_{1},e_{2},e_{3})$ does not directly capture all three-variable interactions, stacking additional KarSein interaction layers effectively addresses this limitation. For instance, in the second KarSein layer, $f^{6}(e_{1},e_{2},e_{3})$ undergoes further activation, resulting in more intricate interactions like $f^{18}(e_{1},e_{2},e_{3})$, which encompasses feature interactions among all three variables. This way, the model successfully learns richer multiplicative relationships.
Table 1: Comparison of different CTR methods.
Methods Feature Interaction Feature De-redundancy FLOPs
High-order Adaptive Original Space
DNN (Covington, Adams, and Sargin 2016) $O(H(mD+H\mathcal{T}))$
KAN (Liu et al. 2024) $O(H(mD+H\mathcal{T})(g+\kappa)$
KarSein-implicit $O((mD+H\mathcal{T})(g+\kappa+H))$
AFN+ (Cheng, Shen, and Huang 2020) $O(mDK+KHD+mHD+H^{2}\mathcal{T})$
EulerNet (Tian et al. 2023) $O(mnD^{2}+n^{2}D^{2}\mathcal{T}))$
KarSein-explicit $O((m^{2}+H\mathcal{T})(g+\kappa+H))$
Activation Transformation
We denote the basis functions of B-Spline curves be of grid size $g$ and order $\kappa$ as $\mathbf{N}_{\kappa}=[N_{1,\kappa},N_{2,\kappa},\ldots,N_{g+\kappa,\kappa}]$. For $L$-th layer’s input matrix $\mathbf{X}^{L-1}\in\mathbb{R}^{H_{L-1}\times D}$. We first activate each row of $\mathbf{X}^{L-1}$ on $g+\kappa$ basis functions, the process is denoted as $\mathbf{X}^{L-1}_{\text{basis}}=\mathbf{N}_{\kappa}(\mathbf{X^{L-1}})\in\mathbb{R}^{H_{L-1}\times D\times(g+\kappa)}$. Then we define learnable weight matrices $\mathbf{C}^{L-1}\in\mathbb{R}^{H_{L-1}\times 1\times(g+\kappa)}$. Then we process the activation transformation for $\mathbf{X}^{L-1}$ using the following formulation:
$\mathbf{X}^{L-1}_{b}=\begin{bmatrix}\mathbf{X}^{L-1}_{\text{basis}}[1,:,:]\mathbf{C}^{L-1}[1,:,:]^{T}\\ \vdots\\ \mathbf{X}^{L-1}_{\text{basis}}[H_{L-1},:,:]\mathbf{C}^{L-1}[H_{L-1},:,:]^{T}\end{bmatrix}$
Linearly Combination
We define the weight matrix $\mathbf{W}^{L-1}_{b}\in\mathbb{R}^{H_{L}\times H_{L-1}}$. To model the feature interactions, we perform a linear combination of the activated embedding vectors, represented as $\mathbf{W}^{L-1}_{b}\mathbf{X}^{L-1}_{b}$. To enhance the expressiveness of the model, we introduce an additional residual connection. Specifically, we apply the $\text{SiLU}(\cdot)$ activation function to $\mathbf{X}$ and define another weight matrix $\mathbf{W}^{L-1}_{s}\in\mathbb{R}^{H_{L}\times H_{L-1}}$ to perform a linear transformation on the activated embeddings, expressed as $\mathbf{W}^{L-1}_{s}\text{SiLU}(\mathbf{X}^{L-1})$. The final output features are then given by the following formulation:
$\mathbf{X}^{L}=\mathbf{W}^{L-1}_{b}\mathbf{X}^{L-1}_{b}+\mathbf{W}^{L-1}_{s}\text{SiLu}(\mathbf{X}^{L-1})$
Integrating Implicit Interactions
We integrate implicit interactions, which is focused on bit-wise level feature interactions. We employ a parallel network architecture that separates the modeling of vector-wise and bit-wise interactions, with both networks sharing the same embedding layer.
For bit-wise interactions, we input a wide concatenated vector $\mathbf{e}\in\mathbb{R}^{mD}$ into the first layer. The total number of stacked KarSein interaction layers is $\mathcal{T}$, the final layer is designed to have a single neuron. The output of the $\mathcal{T}$-th layer is denoted as $\mathbf{e}^{\mathcal{T}}\in\mathbb{R}^{1}$.
CTR Prediction
The final outputs from the KarSein architecture for explicit feature interaction $\mathbf{X}^{\mathcal{T}}$, and the outputs from the KarSein architecture for implicit feature interaction $\mathbf{e}^{\mathcal{T}}$, are combined for binary classification in CTR prediction. The predicted probability, represented by $\hat{y}$, is calculated as follows:
$\hat{y}=\frac{1}{1+\exp(-\mathbf{X}^{\mathcal{T}}\mathbf{W}_{o})}+\frac{1}{1+\exp(-\mathbf{e}^{\mathcal{T}})}$
where $\mathbf{W}_{o}$ represents the regression parameters.
Training with Sparsity
The KAN network exhibits sparsity by applying L1 regularization to the parameters of the activation functions and entropy regularization to the activated values. Our model inherits this feature with enhanced efficiency. Instead of applying L1 regularization to the activation functions’ parameters and entropy regularization to the post-activated values of intermediate input-output features, we incorporate L1 and entropy regularization into the KarSein interaction layer’s linear combination step to eliminate redundant hidden neurons.
Specifically, for the $L$-th layer of the KarSein interaction layer, we apply L1 regularization to $\mathbf{W}^{L-1}_{b}$ and $\mathbf{W}^{L-1}_{s}$ with the regularization parameter $\lambda_{1}$. The L1 regularization term is computed as follows:
$\lambda_{1}||\mathbf{W}^{L-1}_{b}||_{1}+\lambda_{1}||\mathbf{W}^{L-1}_{s}||_{1}$
Next, we compute the entropy regularization term for $\mathbf{W}^{L-1}_{b}$ and $\mathbf{W}^{L-1}_{s}$ with the regularization parameter $\lambda_{2}$. The computation is as follows:
$\lambda_{2}\,\mathcal{H}\left(\frac{\mathbf{W}^{L-1}_{b}}{||\mathbf{W}^{L-1}_{b}||_{1}}\right)+\lambda_{2}\,\mathcal{H}\left(\frac{\mathbf{W}^{L-1}_{s}}{||\mathbf{W}^{L-1}_{s}||_{1}}\right)$
where $\mathcal{H}(\cdot)$ denotes the entropy calculation. The total training objective is given by:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pred}}+\lambda_{1}\sum_{L=1}^{\mathcal{T}}\left(\|\mathbf{W}^{L-1}_{b}\|_{1}+\|\mathbf{W}^{L-1}_{s}\|_{1}\right)+$
$\lambda_{2}\sum_{L=1}^{\mathcal{T}}\left(\mathcal{H}\left(\frac{\mathbf{W}^{L-1}_{b}}{\|\mathbf{W}^{L-1}_{b}\|_{1}}\right)+\mathcal{H}\left(\frac{\mathbf{W}^{L-1}_{s}}{\|\mathbf{W}^{L-1}_{s}\|_{1}}\right)\right)$
Analysis
Let $\mathcal{T}$ is the depth of both KarSein-explicit and KarSein-implicit layers, and $H$ as the number of hidden neurons per layer. Additionally, $K$ represents the number of logarithmic neurons in AFN+ (Cheng, Shen, and Huang 2020), which significantly exceeds $mD$. The parameter $n$ denotes the number of order vectors in EulerNet, is typically set to $m$ in practical applications. Then we present a comprehensive comparative analysis, in terms of floating-point operations (FLOPs), among the proposed KarSein model, KAN, and other state-of-the-art CTR methods, as outlined in Table 1.
We compare DNN, KAN, and KarSein-implicit in the context of implicit feature interaction modeling. DNN demonstrates higher efficiency compared to KAN; however, KAN compensates by requiring significantly fewer hidden neurons $H$ and shallower layers $\mathcal{T}$. The proposed KarSein-implicit method, with parameters $g$ and $\kappa$ much smaller than $H$, achieves computational efficiency comparable to DNN while retaining KAN’s advantage of smaller $H$ and $\mathcal{T}$.
In comparison of AFN+, EulerNet, and KarSein-explicit models which are designed to adaptively learn high-order features, AFN+ and EulerNet involve embedding space transformation. In contrast, KarSein-explicit performs feature interactions directly within the original space. Additionally, KarSein-explicit benefits from KAN’s structural sparsity pruning capability, resulting in more global explainability and feature interactions with less redundancy. Notably, KarSein-explicit exhibits computational complexity independent of $D$, significantly outperforming AFN+ and EulerNet, and even surpassing DNN due to its smaller $\mathcal{T}$ and $H$ (often close to $m\sim m^{2}$).
Table 2: Overall performance and computation cost comparison of different models across three datasets.
Model Criteo Douban Movie MovieLens-1M
AUC($\uparrow$) LogLoss($\downarrow$) Params AUC($\uparrow$) LogLoss($\downarrow$) Params AUC($\uparrow$) LogLoss($\downarrow$) Params
KAN (Liu et al. 2024) 0.8026 0.4472 0.733 M 0.8244 0.3657 0.053 M 0.8273 0.3378 0.083 M
DNN (Covington, Adams, and Sargin 2016) 0.8102 0.4417 0.226 M 0.8305 0.3483 0.021 M 0.8403 0.3287 0.029 M
Wide & Deep (Cheng et al. 2016) 0.8112 0.4409 0.226 M 0.8306 0.3482 0.021 M 0.8410 0.3281 0.029 M
DeepFM (Guo et al. 2017) 0.8131 0.4382 0.113 M 0.8307 0.3483 0.037 M 0.8444 0.3267 0.046 M
xDeepFM (Lian et al. 2018) 0.8132 0.4382 0.656 M 0.8292 0.3516 0.058 M 0.8484 0.3252 0.055 M
DCNV2 (Wang et al. 2021) 0.8135 0.4377 2.244 M 0.8298 0.3502 0.623 M 0.8522 0.3192 0.697 M
AFN+ (Cheng, Shen, and Huang 2020) 0.8129 0.4397 9.982 M 0.8300 0.3493 5.444 M 0.8492 0.3236 5.448 M
FiGNN (Li et al. 2019) 0.8130 0.4385 0.047 M 0.8307 0.3481 0.004 M 0.8475 0.3243 0.006 M
EulerNet (Tian et al. 2023) 0.8129 0.4392 0.303 M 0.8308 0.3480 0.253 M 0.8531 0.3188 0.048 M
KarSein(Ours) 0.8145 0.4372 0.142 M 0.8323 0.3475 0.009 M 0.8555 0.3144 0.018 M
Experiments
In this section, we are to address these research questions: RQ1: How does KarSein model perform compared to other state-of-the-art methods for CTR prediction? RQ2: How do the explicit and implicit components of the KarSein model perform individually in prediction? RQ3: How do different configurations of pairwise multiplication affect the model’s performance? RQ4: What are the learned features in KarSein? RQ5: To what extent can KarSein reduce redundant features and achieve network structural sparsity?
Experiment Setups
Datasets
We conduct experiments on three datasets, including MovieLens 1M, Douban Movie, and Criteo, which have been utilized in previous studies (Cheng, Shen, and Huang 2020; Tian et al. 2023). For each dataset, we randomly split the instances by 8:1:1 for training, validation and test, respectively. Further detailed introduction to our used datasets are in Appendix A.1.
Baseline Methods
We compare our method with three classes of baselines: (1) Methods only have implicit feature interactions, i.e., DNN (Covington, Adams, and Sargin 2016), KAN (Liu et al. 2024), Wide & Deep (Cheng et al. 2016), DCNV2 (Wang et al. 2021). (2) Methods have implicit feature interactions, and explicit feature interactions with predefined order, i.e., DeepFM (Guo et al. 2017), xDeepFM (Lian et al. 2018). (3) Methods have implicit feature interactions, and explicit feature interactions with adaptive order, i.e., AFN+ (Cheng, Shen, and Huang 2020), EulerNet (Tian et al. 2023). (4) Feature interactions built on GNN, i.e., FiGNN (Li et al. 2019). A detailed introduction to these baselines can be found in Appendix A.2
Implementation Details
We utilize Python 3.8, PyTorch 1.13, CUDA 11.6, and a single Quadro RTX 6000 GPU for implementation. For each method, we all conduct extensive grid search over key hyper parameters, and the details can be found in Appendix A.3. For evaluation, we employ AUC and LogLoss as metrics to assess the predictive performance of the models. Additionally, we record the number of parameters, excluding those within the embedding layer, to accurately reflect the computational complexity of each model. For all comparative methods, we run each experiment five times and report the mean results.
Overall Performance (RQ1)
This section provides a comparative analysis of the performance and parameter computation cost (excluding the embedding table parameters) between the proposed KarSein model and existing state-of-the-art baselines for CTR prediction. The experimental results are summarized in Table 2.
In terms of performance, KarSein consistently outperforms all baseline methods across three datasets. This demonstrates the efficacy of our learnable activation functions in capturing high-order feature interactions. Notably, all CTR models achieve an AUC of approximately 0.83 on the Douban dataset. Even advanced models like EulerNet struggle to achieve a 0.001 increase in AUC. However, our approach significantly surpasses these methods, delivering an improvement of 0.002 over the 0.83 benchmark. This enhancement is particularly noteworthy, given that previous studies (Cheng et al. 2016; Guo et al. 2017; Wang et al. 2021) have established that an AUC increase or Logloss reduction at the 0.001 level is statistically significant.
Regarding computational efficiency, the parameter computation cost associated with the KarSein model is remarkably lower than that of several SOTA methods, even outperforming traditional DNN models. This finding aligns with our previous FLOPs analysis.
Ablation Study (RQ2)
We investigate the contributions of explicit and implicit feature interactions in the KarSein model by isolating and evaluating each type independently. Our analysis involves decomposing the default ensemble KarSein model, which combines both interaction types, to assess the impact of using only explicit or only implicit interactions. We compare these models in terms of CTR prediction performance, model size, and training time. The results, detailed in Table 3, show that the KarSein model with only explicit interactions achieves similar AUC performance to the full ensemble model but with smaller model size and reduced training time. Conversely, the model with only implicit interactions shows lower AUC performance and a significant increase in model parameters. Our findings highlight that while combining both interaction types yields superior AUC performance, explicit interactions offer advantages in efficiency and parameter size. Implicit interactions, although adding considerable model complexity, have minimal impact on training time. This underscores the complementary strengths of each interaction mechanism.
Table 3: Impact of Feature Interaction Mechanisms on KarSein Model: Performance, Parameter Size, and Training Time.
Model Criteo Douban Movie MovieLens-1M
AUC Params Time×Epochs AUC Params Time×Epochs AUC Params Time×Epochs
KarSein-explicit 0.8142 0.051 K 30m × 3 0.8319 0.269 K 14.4s × 3 0.8542 0.812 K 8.4s × 3
KarSein-implicit 0.8125 0.091 M 23m × 4 0.8311 0.008 M 13.2s × 2 0.8535 0.017 M 5.7s × 10
KarSein 0.8145 0.092 M 40m × 3 0.8323 0.009 M 19.5s × 2 0.8555 0.018 M 9.9s × 3
Robustness Study (RQ3)
We empirically demonstrate that performing pairwise multiplication in only the first two KarSein interaction layers is sufficient for the model to learn multiplicative relationships effectively. Figure 3 illustrates the impact of applying pairwise multiplication across different layers on KarSein-explicit’s performance using the MovieLens-1M and Douban datasets. When pairwise multiplication is omitted (Layer Index set to None), the model’s AUC performance is notably poor. Introducing this step in just the first layer leads to significant performance gains, especially on the MovieLens-1M dataset, which benefits from its six basic field features. In contrast, the Douban dataset, with only two features, shows less improvement.
These results also confirm our earlier statement that the KAN network has difficulty learning multiplicative relationships on its own, limiting its CTR prediction effectiveness. Furthermore, applying pairwise multiplication exclusively in the first two layers achieves the highest AUC scores across both datasets. Deviations from this setup—such as using only one layer or extending beyond the first two layers—result in decreased performance.
Figure 3: Impact of Pairwise Multiplication in Different Configurations on AUC Performance of the KarSein-explicit
Explanation Study (RQ4)
Our further exploration of the KarSein model focuses on how learnable activation functions transform low-order input features into output features. We visualized these activation functions across various layers and used third-degree polynomials for symbolic regression. Most activation functions were well-approximated by cubic polynomials, as shown in the first row of Figure 4. This indicates that KarSein’s activation functions effectively elevate low-order features to capture high-order interactions.
We also observed some activation functions displaying oscillatory and irregular patterns, depicted in the second column of Figure 4. This behavior underscores the effectiveness of B-Spline activation functions, which offer the flexibility to model such complex patterns. This capability to model high-order and intricate features significantly contributes to KarSein’s state-of-the-art performance.
Pruning Redundant Features (RQ5)
We assess the level of sparsity and feature redundancy reduction in the optimized KarSein model. Specifically, we trained the KarSein model on the MovieLens-1M dataset and achieved an AUC of 0.8555. We use heat maps to visualize connections from activated inputs to outputs across each layer of the model, and the detailed results are in Appendix C. We find that certain input features do no contribute to any output features (values $\leq 0.01$). These features are deemed non-essential and can be masked. In the KarSein-explicit component, 66% of input features in the first layer are redundant. For the KarSein-implicit component, 83%, 87%, and 44% of input features in the first, second, and third layers, respectively, are redundant. We then masked these redundant features and continued training the KarSein model for an additional 3 epochs until convergence, resulting in an AUC of 0.8533. This demonstrates the model’s ability to remove redundant features without greatly compromising performance. This characteristic surpasses many pioneer CTR methods that only can provide contextual feature importance, achieving feature ”de-redundancy.” The resulting network structural sparsity may can be leveraged to accelerate inference in recommendation systems.
Figure 4: Visualization of several representative learnable activation functions in KarSein with $\kappa=3$, $g=10$
Related Works
Adaptive Order Feature Interaction Learning
Recent advancements in CTR prediction aim to surpass the limits of predefined interaction orders and feature combinations, improving predictive performance while reducing the computational burden of enumerating high-order interactions. Notable representatives are AFN (Cheng, Shen, and Huang 2020) and EulerNet (Tian et al. 2023). Both methods shift the multiplication-based feature interactions traditionally performed in the original embedding space to an alternative space where feature combinations are achieved through linear operations before being mapped back to the original space. Specifically, AFN employs logarithmic transformation (Hines 1996), whereas EulerNet utilizes Euler’s formula. However, these methods introduce challenges related to numerical stability issue, potential loss of feature relationships, and additionally computational overhead. In response, we propose the KarSein method, which directly activates features within the original space to adaptive orders using a learnable activation function. Our approach is much more simpler and effective, also has intuitive symbolic regression explanations.
Feature Importance Learning
Techniques, such as AFM (Xiao et al. 2017), AutoInt (Weiping et al. 2018), FiGNN (Li et al. 2019), and FiBiNET (Huang, Zhang, and Zhang 2019), leverage self-attention (Vaswani 2017) mechanisms or SENet (Hu, Shen, and Sun 2018) to provide contextual explanations for the prediction process. However, the explainability afforded by these methods is inherently local, limiting their ability to identify universally redundant features across all samples. In contrast, the global interpretability of the KarSein method allows for the pruning of unnecessary network structures by eliminating redundant feature learning, resulting in a sparser network architecture. This sparsity holds significant potential for efficient inference.
Conclusion and Future Work
In this paper, we introduced KarSein, a Kolmogorov-Arnold Represented Sparse Efficient Interaction Network, to address the limitations of traditional CTR prediction methods. KarSein enhances predictive accuracy and reduces computational costs through a novel architecture that uses learnable activation functions for modeling high order features from low order features efficiently. It employs enforced pairwise multiplication, guiding KarSein’s symbolic regression incorporates multiplicative relationships. Our extensive experiments demonstrate KarSein’s superior performance, maintaining global interpretability and a sparse network structure, thus positioning it as a highly efficient method for CTR prediction.
Appendics
Appendix A: Experiment Setups
Appendix A.1: Datasets
We conduct experiments on three datasets, including MovieLens 1M, Douban Movie, and Criteo.
MovieLens 1M : MovieLens 1M dataset is a widely used benchmark dataset in the field of recommender systems. It consists of 1 million movie ratings provided by 6,000 users on 4,000 movies. Each rating ranges from 1 to 5 for rating prediction. For CTR prediction, the ratings of 1 and 2 are normalized to be 0, and ratings of 4 and 5 to be 1.
Douban Movie: The Douban Movie dataset consists of 1 million movie ratings, which are collected from the Douban website and ranged between 1 to 5 for rating prediction. The dataset includes data from 10,000 users and 10,000 movies spanning the years 2008 to 2019. For CTR prediction, the ratings of 1 and 2 are normalized to be 0, and ratings of 4 and 5 to be 1.
Criteo Dataset: The Criteo dataset comprises user logs collected over a period of 7 days. It contains 45 million examples and 39 features, including 26 categorical feature fields and 13 numerical feature fields. We discretize each numerical value using a logarithmic discretization method.
Appendix A.2: Baseline Methods
We compare KarSein with state-of-the-art methods in CTR prediction task, including:
  • DNN (Covington, Adams, and Sargin 2016) is a straightforward model based on deep stacked MLP architecture , which applies a fully-connected network after the concatenation of feature embeddings for CTR prediction.
  • KAN (Liu et al. 2024) offers a promising alternative to MLPs. In our approach, we replace MLPs with KAN, allowing concatenated feature embeddings to be processed through KAN for CTR prediction.
  • Wide & Deep (Cheng et al. 2016) combines a linear model for memorization of feature interactions with a DNN for generalization to capture high-order interactions.
  • DCNV2 (Wang et al. 2021) uses the kernel product of concatenated feature vectors to model high-order interactions and integrates an DNN for implicit interactions.
  • DeepFM (Guo et al. 2017) combines FM to capture second-order interactions with DNN to model high-order interactions.
  • xDeepFM (Lian et al. 2018) encodes high-order interactions into multiple feature maps and combines them with an DNN to model implicit interactions.
  • AFN+ (Cheng, Shen, and Huang 2020) AFN encodes features into a logarithmic space to adaptively learn arbitrary-order feature interactions, AFN+ additionally use an DNN for implicit interactions, further improving the model capacity.
  • EulerNet (Tian et al. 2023) leverages Euler’s formula to transform features into the complex vector space, allowing for the efficient modeling high order feature interactions linearly.
  • FiGNN (Li et al. 2019) represents features as a fully-connected graph and uses gated Graph Neural Networks (GNNs) to model high-order feature interactions.
Appendix A.3: Hyper Parameter Setting
For the sake of fair comparison, the embedding size is uniformly set to 16 across all methods. For each method, we all conduct grid search over several general hyper parameters to optimize performance. The learning rate is selected from $\{1e^{-3},2e^{-3},3e^{-3},1e^{-4}\}$, while the batch size is varied across $\{512,1024,4096\}$. Many methods incorporate the DNN components, we experimente with hidden layer configurations of $\{400-400-400,128-128-128,256-256-256\}$, with a dropout rate of 0.1. The regularization weight penalty is chosen from $\{1e^{-5},1e^{-6}\}$.
We conduct further grid search for tuning the unique hyper parameters associated within each method. In the xDeepFM model, the Compressed Interaction Network (CIN) depth is varied among ${1,2,3}$, with the hidden size chosen from $\{100,200\}$. For DCNV2 model, the low-rank space size is set to 128, while the CrossNet depth is tested across $\{3,4,5\}$. For FiGNN model, the number of attention heads is set to 2, and the graph interaction steps are chosen from $\{2,3,4\}$. For AFN, the number of logarithmic neurons is selected from $\{400,800,1000,1500\}$. For EulerNet, the number of Euler interaction layers is selected from $\{1,2,3\}$, and the number of order vectors is fixed at 30.
For the implicit components of our proposed KarSein model, we choose hidden layers from $\{32-32,64-64,64-32\}$ for all datasets. For the explicit components, we choose the hidden layers from $\{50-50,26-26,16-16\}$ for the Criteo dataset, and $\{4-4,6-6,8-8\}$ for the Douban and MovieLens-1M datasets. Regularization parameters $\lambda_{1}$ and $\lambda_{2}$ are selected from $\{1e^{-2},1e^{-3}\}$ and $\{1e^{-4},1e^{-5}\}$ respectively. The $\kappa$ and $g$ are set from $\{(\kappa=1,g=3),(\kappa=2,g=5),(\kappa=3,g=10)\}$.
Table 4: Steps required to achieve RMSE $\leq 0.05$ for three different KAN settings in fitting $a^{2}$, $b^{2}$, and $ab$. The symbol \ indicates the configuration is unable to reach the specified RMSE.
Setting ID KAN Layers Regularization Steps to Converge
$a^{2}$ $b^{2}$ $ab$ $a^{2}$ $b^{2}$ $ab$
1 2 - 1 2 - 1 2 - 2 - 1 0.01 600 600 1200
2 2 - 4 - 1 2 - 4 - 1 2 - 2 - 4 - 1 0.01 350 350 \
3 2 - 4 - 1 2 - 4 - 1 2 - 2 - 4 - 1 0.00 250 250 320
Appendix B: KAN for CTR Prediction
Appendix B.1: Exploration 1
We are to exploration whether KAN structure can spontaneously learn multiplicative interactions. Given two basic feature $a$ and $b$. We pre-define the idealized KAN structures $2-1$, $2-1$, and $2-2-1$ for learning these interactions: $f(a,b)=a^{2}$, $f(a,b)=b^{2}$, and $f(a,b)=ab$, respectively. Beyond this idealized framework, our investigation extends to two additional network initialization settings, one incorporating regularization and the other excluding it. Comprehensive details of these configurations are delineated in Table 4.
Our primary objective is to model second-order feature interactions to achieve a RMSE of $\leq 0.05$. We employ the adam optimizer with a learning rate of 0.001 in all configurations. We document the number of optimization steps required to meet this criterion. Additionally, we present visualizations of the model structures upon reaching the target, as shown in Figure 5. From the experimental results, we observe the following:
  • Under predefined ideal structures, the KAN not only accurately fits the feature interactions but also effectively performs symbolic regression, successfully learning the multiplicative relationship.
  • In non-ideal structures with regularization, the KAN fails to automatically perform symbolic regression and learn the multiplicative relationship. Additionally, it also fails to fit $ab$ because of setting to high regularization making the structural too sparse.
  • In non-ideal structures without regularization, the KAN aptly fits all second-order feature expressions but does not succeed in symbolic regression.
Figure 5: Visualization of KAN for fitting simple second-order feature interactions across three different settings.
Conclusion: The KAN demonstrates exceptional capability in fitting feature interactions. When set with appropriate regularization penalty, albeit parameter-sensitive, KAN can achieve excellent fitting performance while maintaining structural sparsity. Besides, KAN is capable of learning multiplicative cross-features only under an idealized initial network structure, which is often requiring prior knowledge. However, in more general scenarios without pre-defined structures and without manual intervention to prune the network during training, KAN cannot reliably and spontaneously perform symbolic regression for learning higher-order multiplicative cross-features. This observation underscores a regrettable reality: KAN, akin to DNN, remains intrinsically limited in its ability to autonomously learn multiplicative feature-based interactions. Consequently, relying solely on KAN’s spontaneously learning may not yield optimal CTR prediction results.
Appendix B.2: Exploration 2
We explore the performance of directly applying KAN in CTR, this study is conducted on the MovieLens-1M dataset. We set the regularization parameter to 0.01, the learning rate to 0.01, and the batch size to 512. The grid size is 3, the B-Spline degree is 1, the embedding dimension $D$ is 16, and $m$ is 6. We train the KAN model until convergence, achieving an AUC score of 0.8273 (A DNN model with the same layer and neuron configuration as KAN can achieve an AUC of 0.8403). Subsequently, we pruned nodes with small incoming or outgoing connections (weights $\leq 0.003$), resulting in the structure depicted in Figure 6. From the figure, we can observe that the KAN network is pruned from the initial $96-64-64-1$ neurons down to $96-1-1-1$ neurons.
Figure 6: Visualization of KAN for CTR prediction.
Conclusion: The vanilla application of KAN for CTR prediction yields suboptimal results. Additionally, we find that using KAN for the CTR task is parameter-redundant; a single activation function per feature may be sufficient, eliminating the cost for multiple activation functions on the same feature.
Figure 7: Visualization of input-to-output feature connections across layers in $\text{KarSein}_{\text{bit-wise}}$ model: Each row represents an input feature and each column represents an output feature. The color intensity indicates the magnitude of the connection weights.
Appendix C: Pruning Redundant Features
We use heat maps to visualize connections from activated inputs to outputs across each layer of the model, as shown in Figure 7. The heat maps reveal that certain input features contribute minimally to all output features (values $\leq 0.01$ ), indicated by large blank areas. The KarSein-explicit component exhibits fewer redundant features compared to the KarSein-implicit component. This phenomenon aligns with our conclusion that the KarSein-explicit component is the primary contributor to overall performance, with the KarSein-implicit component serving a complementary role.
Appendix D: Feature Interaction Types
Learning feature interactions are crucial for CTR prediction. Current CTR methods’ approach in modeling feature interactions can be broadly categorized into two primary paradigms: implicit and explicit. The implicit paradigm primarily employs deep learning techniques, such as Deep Neural Networks (DNNs) (Covington, Adams, and Sargin 2016), to automatically model latent feature interactions, capturing complex and unknown patterns within the data. Methods in this category rely on the Universal Approximation Theorem. The explicit paradigm, on the other hand, explicitly enumerates feature combinations which are often constructed in a way of inner product of basic field features’ representing vectors. They can often be mathematically represented as higher-order polynomials of the basic field features and are commonly implemented using factorization-based architectures (Pan et al. 2018; Guo et al. 2017; Lian et al. 2018; Naumov et al. 2019; He and Chua 2017). Previous studies verified that the multiplicative product relationships are important for CTR prediction, and are challenging for DNN to spontaneously learn (Weiping et al. 2018; Rendle et al. 2020). Thus, the component of modeling explicit feature interactions is the core for CTR prediction, and current CTR models typically leverage implicit feature interactions as a supplementary signal to the explicit feature interaction component, resulting in an ensemble architecture (Cheng et al. 2016; Naumov et al. 2019; Guo et al. 2017; Lian et al. 2018; Yu et al. 2020; Wang et al. 2021).

### v4

@@
-Beyond KAN: Introducing KarSein for Adaptive High-Order Feature Interaction Modeling in CTR Prediction
+CTR-KAN: KAN for Adaptive High-Order Feature Interaction Modeling
 Abstract
-Modeling feature interactions is crucial for click-through rate (CTR) prediction, particularly when it comes to high-order explicit interactions.
-Traditional methods struggle with this task because they often pre-define a maximum interaction order, which relies heavily on prior knowledge and can limit the model’s effectiveness.
-Additionally, modeling high-order interactions typically leads to increased computational costs.
-Therefore, the challenge lies in adaptively modeling high-order feature interactions while maintaining efficiency.
-To address this issue, we introduce Kolmogorov-Arnold Represented Sparse Efficient Interaction Network (KarSein), designed to optimize both predictive accuracy and computational efficiency.
-We firstly identify limitations of directly applying Kolmogorov-Arnold Networks (KAN) to CTR, and then introduce KarSein to overcome these issues.
-It features a novel architecture that reduces the computational costs of KAN and supports embedding vectors as feature inputs.
-Additionally, KarSein employs guided symbolic regression to address the challenge of KAN in spontaneously learning multiplicative relationships.
-Extensive experiments demonstrate KarSein’s superior performance, achieving significant predictive accuracy with minimal computational overhead.
-Furthermore, KarSein maintains strong global explainality while enabling the removal of redundant features, resulting in a sparse network structure.
-These advantages also position KarSein as a promising method for efficient inference  [^1] https://github.com/Ancientshi/KarSein.
+Modeling high-order feature interactions is critical for click-through rate (CTR) prediction, yet traditional approaches often face challenges in balancing predictive accuracy and computational efficiency.
+These methods typically rely on pre-defined interaction orders, which limit flexibility and require extensive prior knowledge.
+Moreover, explicitly modeling high-order interactions can lead to significant computational overhead.
+To tackle these challenges, we propose CTR-KAN, an adaptive framework for efficient high-order feature interaction modeling.
+CTR-KAN builds upon the Kolmogorov-Arnold Network (KAN) paradigm, addressing its limitations in CTR prediction tasks.
+Specifically, we introduce key enhancements, including a lightweight architecture that reduces the computational complexity of KAN and supports embedding-based feature representations.
+Additionally, CTR-KAN integrates guided symbolic regression to effectively capture multiplicative relationships, a known challenge in standard KAN implementations.
+Extensive experiments demonstrate that CTR-KAN achieves state-of-the-art predictive accuracy with significantly lower computational costs.
+Its sparse network structure also facilitates feature pruning and enhances global interpretability, making CTR-KAN a powerful tool for efficient inference in real-world CTR prediction scenarios.
+The source code is available at  [^1] https://github.com/Ancientshi/KarSein.
 Introduction

### v5

@@
-CTR-KAN: KAN for Adaptive High-Order Feature Interaction Modeling
+Beyond KAN: Introducing KarSein for Adaptive High-Order Feature Interaction Modeling in CTR Prediction
 Abstract
-Modeling high-order feature interactions is critical for click-through rate (CTR) prediction, yet traditional approaches often face challenges in balancing predictive accuracy and computational efficiency.
-These methods typically rely on pre-defined interaction orders, which limit flexibility and require extensive prior knowledge.
-Moreover, explicitly modeling high-order interactions can lead to significant computational overhead.
-To tackle these challenges, we propose CTR-KAN, an adaptive framework for efficient high-order feature interaction modeling.
-CTR-KAN builds upon the Kolmogorov-Arnold Network (KAN) paradigm, addressing its limitations in CTR prediction tasks.
-Specifically, we introduce key enhancements, including a lightweight architecture that reduces the computational complexity of KAN and supports embedding-based feature representations.
-Additionally, CTR-KAN integrates guided symbolic regression to effectively capture multiplicative relationships, a known challenge in standard KAN implementations.
-Extensive experiments demonstrate that CTR-KAN achieves state-of-the-art predictive accuracy with significantly lower computational costs.
-Its sparse network structure also facilitates feature pruning and enhances global interpretability, making CTR-KAN a powerful tool for efficient inference in real-world CTR prediction scenarios.
-The source code is available at  [^1] https://github.com/Ancientshi/KarSein.
+Modeling high-order feature interactions is crucial for click-through rate (CTR) prediction, and traditional approaches often predefine a maximum interaction order and rely on exhaustive enumeration of feature combinations up to this predefined order.
+This framework heavily relies on prior domain knowledge to define interaction scope and entails high computational costs from enumeration.
+Conventional CTR models face a trade-off between improving representation through complex high-order feature interactions and reducing computational inefficiencies associated with these processes.
+To address this dual challenge, this study introduces the Kolmogorov-Arnold Represented Sparse Efficient Interaction Network (KarSein).
+Drawing inspiration from the learnable activation mechanism in the Kolmogorov-Arnold Network (KAN), KarSein leverages this mechanism to adaptively transform low-order basic features into high-order feature interactions, offering a novel approach to feature interaction modeling.
+KarSein extends the capabilities of KAN by introducing a more efficient architecture that significantly reduces computational costs while accommodating two-dimensional embedding vectors as feature inputs.
+Furthermore, it overcomes the limitation of KAN’s its inability to spontaneously capture multiplicative relationships among features.
+Extensive experiments highlight the superiority of KarSein, demonstrating its ability to surpass not only the vanilla implementation of KAN in CTR prediction tasks but also other baseline methods.
+Remarkably, KarSein achieves exceptional predictive accuracy while maintaining a highly compact parameter size and minimal computational overhead.
+Moreover, KarSein retains the key advantages of KAN, such as strong interpretability and structural sparsity.
+As the first attempt to apply KAN in the CTR domain, this work introduces KarSein as a novel solution for modeling complex feature interactions, underscoring its transformative potential in advancing CTR prediction task.
+The code is available at  [^1] https://github.com/Ancientshi/KarSein.
 Introduction